OpenStudy (anonymous):

Determine if the expression is a perfect square and if it is, factor it. a^2 – 8a + 16 A. (a + 4)(a – 4) B. (a – 4)(a – 4) C. not a perfect square D. (4a + 1)2

OpenStudy (shamil98):

ax^2 + bx + c = 0 c is positive and b is negative negative x negative = positive so your answer based on this deduction would be?..

OpenStudy (anonymous):

A right @shamil98

OpenStudy (shamil98):

-4 + (-4) = -8 right? and -4 x - 4= 16 think about it.

OpenStudy (anonymous):

Oh ok im totally not guessing I thought about it and Its B @shamil98

OpenStudy (anonymous):

Right?

OpenStudy (shamil98):

Yes.